OpenStudy (anonymous):

The sum of the measure of angle X and angle Y is 90. If the measure of angle X is 30 less than twice the measure of angle Y, what is the measure of angle X?

OpenStudy (anonymous):

x + y = 90 x = 2y - 30 solve the eqaution

OpenStudy (anonymous):

multiply 1st eqaution x+y =90 by 2 so it would be 2x + 2y = 180 then add the second equation to the 1st i.e (2x + 2y = 180 ) + (x-2y = -30) so 3x= 150 x = 50

OpenStudy (anonymous):

hope u understood

OpenStudy (anonymous):

just like wat they said, but instead u can just do this,